An element with mass 630 grams decays by 18.8% per minute. How much of the element is remaining after 7 minutes, to the nearest 10th of a gram?



1
Expert's answer
2022-02-02T16:19:18-0500

Rate, "r=18.8\%"

Time, "t=7"

Initial mass, "i=630"

To calculate the remaining element, we apply the formula, "i(1-{r\over100})^t=630(1-0.188)^7=630\times 0.2328=146.6"

Therefore, the remaining element after 7 minutes is 146.6 grams (to the nearest 10th of a gram).